sorry for the last incomplete question the original one is:
A non volatile organic compound is used to make two solutions A,B .Solution A contains 5 gm of Z dissolved in 100 gm of water and solution B contains 2.31 gm of Z in 100 gm of benzene .Solution A has a vapour pressure of 754.5 mm Hg at the normal boiling point of water and solution B has the same vapour pressure at the normal boiling point of benzene.calculate the molar mass of Z in the solutions A and B and account for the difference

Grade:12

1 Answers

Aman Bansal
11 years ago

Dear Dev

 Given data
  Mass of water = 100 g
  The mass of solute = 5 g
  Number of moles of water = mass / molar mass
                                          = 100g/18g/mol
                                       n2= 5.56 mol
  The vapor pressure of water at 100 deg C, P0 = 760mmHg
    The vapor pressure of solution at 100 deg C= 754.5 mmHg
   We know that
  Accroding to Raoults Law
    The vapor pressure of solution  =mole fraction of water *P0
               Molefraction of water = 754.5 mmHg / 760 mmHg
                                                  =0.9927
   Mole fraction of solute = 1-0.9927
                                      =0.00723
             Molefraction of solute =n1/(n1+n2
                    ∴ n1/(n1+n2) =0.00723
                        (n1+n2) /n1 = 1/0.00723
                          (n1+5.56mol) /n1  = 138.3
                                                 n1 = 0.04 mol
     Molar mass of solute Z = mass /moles
                                        = 5.0 g / 0.04 mol
                                        = 125 g/mol
 
CaseII
  Given data
  Mass of benzene = 100 g
  The mass of solute = 2.31 g
  Number of moles of benzene = mass / molar mass
                                            = 100g/78 g/mol
                                        n2 = 1.28 mol
  The vapor pressure of benzene at 100 deg C, P0 =760 mmHg
    The vapor pressure of solution at 100 deg C= 754.5 mmHg
   We know that
  Accroding to Raoults Law
    The vapor pressure of solution  =mole fraction of benzene *P0
               Molefraction of water = 754.5 mmHg / 760 mmHg
                                                  =0.9927
   Mole fraction of solute = 1-0.9927
                                      =0.00723
             Molefraction of solute =n1/(n1+n2
                    ∴ n1/(n1+n2) =0.00723
                        (n1+n2) /n1 = 1/0.00723
                          (n1+1.28mol) /n1  = 138.3
                                                 n1 = 0.00932 mol
     Molar mass of solute Z = mass /moles
                                        = 2.31 g / 0.00932 mol
                                        = 247.85 g/mol
      Molar mass of Z in case II isalmost double to the molar mass of Z in case I that means thesolute dimerizes in bezene.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert